Systems Staff Meeting Notes 10/9/00

Systems Staff Meeting Notes
October 9, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. FSU lost!
  2. Winston asks if he can get rid of his beeper. Not!
  3. The Windows 2000 class was cancelled this past Saturday. It has been rescheduled for this Saturday, October 14.
  4. Bobbie's daughter got 2nd in her class, Runner Up High Point, in barrel racing (horseback riding).
  5. Items for Systems - carry forward money - server room: replacement servers, UPS, security & fiber wiring. Bill mentioned to Will that he would like to buy cluster servers. That would be a total of 6 servers - 2 for the SQL, 2 for Email and 2 for the Web server. These run about $9000 each, $5000 for racks. Martha would like more specifics on server room items. Will check with Bill on this. Will notes that we would also need a back-up tape drive - DLT type - which would hold 7 tapes & requiring changing the tape only once per week. No word on the list of tapes to be eliminated from the collection. Need to contact Dolores again.
    New furniture to be requested for Debra/student area.
    Phone & network connections to be requested for the repair tech work area.
    Conference room furniture.
    New digital camera to be made available for checkout.
  6. Summary of Martha's LMS meeting at UCF - The document with requirements will be sent out to vendors this week. Response expected by December 1st. Decision should be made by the end of this fiscal year. The proposal is to implement it library-by-library. Will do the smallest library first (Gulf Coast University) and then UF (because of acceptance testing). Access method is client/server based with Web based technical side.
Review Minutes of 10/02/00
  1. Liaisons #1 - Training Plans - Martha notes that at the NERDC I&R meeting they were discussing NetG courses & she was thinking about the possibility of the liaisons using them to learn some basics. They have pre-tests that you can take - maybe it could help with network & hardware information. We could talk to NERDC & get a subset of the courses set up on our server, allowing us to have Admin access to monitor it. It could be used as a pre-assessment for liaison skill levels. Generally, the courses are a more in depth course than what we need. Both Ying & Cynthia have had some experience with NetG courses - they say that sometimes you'll have the concept but you must know the exact answer. It is very, very literal.
  2. No feedback on the CIRCA course attended by some liaisons last week. Will poll the first group for their reaction.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 20 Dells arrived last week. 15" monitors were ordered & received, but they had to be returned for 17" ones. Hopefully they will arrive within a week. The CPU's & speakers are all in the vault, except one CPU that was issued to Rob for the Security System. There was a bit of controversy with the late-hour delivery. Let it be known that with an inside delivery, drivers must call 24 hours in advance. We have the right to refuse acceptance otherwise. Once the new monitors arrive & we start getting these machines out, we can begin to work on another order.
  2. DLC machines arrived during regular work hours. They've been checked in and are currently located in the outer vault. We'll let DLC know that we've received them, although they may want to wait until they've relocated to the 2nd floor before picking up the machines.
  3. LAC has returned a machine, one to be replaced with a new Dell. Documents will be returning WSID 1116, rather than WSID 1024.
  4. Barcode Readers - returned ones to the vendor from MSL. They are going to replace them with six contact-type scanners. We received six more long-range ones, but they were returned as the order had already been cancelled. Resource Services returned three long-range scanners on Friday.
  5. Laptops - three in the vault - one will be for Reference Chat (Jana or Colleen are the contact people). The other two will be used to replace the Toshibas that we use for checkout.
  6. Eleven more laptops are on order. We may use two of them to upgrade the remaining two Toshibas available for checkout. The others will be used for training purposes.
  7. We will buy a new server for Ref Chat/Net Agent. A dedicated server is required.
  8. Fiber switch order has been forwarded to Barbara Oliver.
  9. Betty Mitchell called to say that she received a bill from Best Buy for a Visor order that was cancelled. We think we sent those Visors back. Will check the paperwork.
  10. A new CISCO switch installed in 123 West on Thursday was dead this morning. Bobbie put our last copper switch out in order to replace it.
Liaisons
  1. Training Session - preparation is in progress. Will try to limit it to what they really need to know. Geoffrey is currently trimming down & compiling the information.
Networking
  1. Public Printing - no change.
  2. Proxy server - keeps going up and down. Will is trying to learn Linux in order to build a new server.
  3. Cataloger's Desktop - not installed yet.
  4. Rewire for DLC & Resource Services - fiber in North tower terminated on the 3rd & 4th floor.
  5. NT9 needs to be relocated or a rack configuration needs to be built for it.
  6. South tower wiring - still have three pulls to be terminated in room 310. Need to remove punch down & HUB in Susan Lupi's office.
  7. Suzy's network maps - still in progress; she scanned in floorplans.
  8. Documents - they moved last week - to places with no electric, network or phone outlets. Requests for such moves should come through Facilities. Need to recheck current wiring encumbrances - David Hickey, Leilani & Gary are all moving. The rooms where drops have been encumbered may not be necessary now. It'd be nice if such moves were planned out and Systems included in the notification.
Software Projects Status
  1. Film Log has been wrapped up on our end. Preservation is checking it out. Will be a couple of weeks before the data transfer.
  2. IR database - Bobbie has requested a couple of changes - Hilton will work on them. Cynthia is working on a read-only interface for Facilities to be able to access certain items in the database.
  3. Michele, Donna & Winston met to discuss the Title ID Conversion. Winston was able to get the macros to run - no documentation on the interface. He got most everything converted over - need another macro to check it, commanding Host Explorer. Most order numbers are fixed.
  4. DLC - Ying is scheduled to meet with Erich tomorrow. She will call down the last version of the code today & make it ready for demo tomorrow.
  5. The Web version of PAULA is up and running.
DP Coordinator
  1. Need to check with Leilani to see if the patron she was dealing with is truly interested in purchasing the NSUR tape.
Calendar
  1. Winston out Oct. 16, 19 & 20.
  2. Ying out Oct. 18-20.
  3. Will and Hilton both out Nov. 20-22 & 27.


Back to the Systems Homepage

Last updated October 16, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu